What was the incidence rate of CHD mortality among women who were current smokers, per 1000 years of follow up? (Use Table 1 to perform this calculation)

TABLE 1. Relative rlsks of coronary heart dlsease (CHD) mortality accordlng to major cardlovascular rlak factore: Flnnlsh Mobllo Clnlc Study, 1988-1984 Men (n - 2,748) Person Relatveconfiance Women (n 2,385) No.of Pei901* Rolatve conidenc No, ol 95% 95% CHO years ๙ cases lolow-up CHD years of Inlervalcases foliow-up dak risk Interval Age (years)t 30-39 40-49 50-59 60-69 1814,943 48 12,560 69 7,677 2 12,627 10 10,972 1.3-26.3 7,63818.5 3.9-70.6 3,86642.6 0.1-179.3 3.21.8-5.5 7,5 4.4-12.5 20 5.8 3,16913.47.8-22.9 28 Smoklng Nonsmoker Current smoker 72 19,090 114 19,259 41 29,827 17 1.8 1.3-2.4 5,114 4.0 2.3-7.2 Body mass index terlle (kg/m2) s24.0 24.1--26.7 >28.7 68 2,585 48 12,925 70 12,823 1311,796 1811,707 27 11,518 0.7 0.5-1.0 0.9 1.0 0,5-1.0 0.6-1.9 Cholesteroi tertile (mgdi) S237 238-280 >280 36 13,079 5712,946 93 12,324 1.40.8-2.2 2.41.7-3.8 15 11,818 11 11,004 32 11,379 0.2-1.1 0.6-1.9 0.5 Hyportensknt Normotensive Borderllne Mild Dellnite 5317,539 113 18,385 986 1,455 1.61.1-2.2 0.6 0.2-2.1 1.01,1-3.4 816,081 3214,040 1,298 3,743 1.90.04.3 2.2 2.1 0.9-5.4 0.6-7.3 Ago adjusted f No adjustment. t Dofnte hypertenslon: systollc blood pressure (SBP) of 170 mmHg an diastolic blood preasuro (DBP) ot 100 mmHg or anthypertensive medicatilon; mld: S0P ol z160 mmHg and DBP ol 295 mmHg, but nol dofnod as dofinite; nomotensve: SBP of <140 mmHg and DBP of «00 mmHg; borderlne: pereons not definod aa doflntoly or mldly hypertongivo or nomotensive
Answer #1

Number of nonsmoking women sampled = 29,827 person-years

Number of CHD cases = 41

Incidence rate of CHD mortality among nonsmoking women, per 1000 years of follow up

= (41 / 29,827) * 1000 = 1.374593

Number of nonsmoking women sampled = 5114 person-years

Number of CHD cases = 17

Incidence rate of CHD mortality among women who were current smokers, per 1000 years of follow up

= (17 / 5114) * 1000 = 3.324208